OpenStudy (anonymous):

A particle, initially at rest, moves along the x-axis such that its acceleration at time t>0 is given by a(t)=cos(t). At the time t=0, its position is x=3. How do I find the velocity and position functions for the particle? I tried integrating the equation but had to integrate a sin(t) + C, something I don't know how to do. Please help! Thanks :D

OpenStudy (anonymous):

you got the main part already. what you missed is that initially it is at rest, so this means \[v(0)=0\]

OpenStudy (anonymous):

so since \[v(t)=\sin(t)+C\] and \[v(0)=0\] you have \[v(0)=\sin(0)+C=0\] and therefore \[C=0\]

OpenStudy (anonymous):

therefore \[v(t)=\sin(t)\] exactly. now take the antiderivative again. you get \[S(t)=-\cos(t)+C\] and this time you know \[S(0)=-\cos(0)+C=3\] so \[-1+C=3\] and therefore \[C=4\]